# RAGDiffusion: Faithful Cloth Generation via External Knowledge Assimilation

### v1

@@ 4.1 Experimental setup
Datasets. We collected 65,131 pairs of (in-the-wild upper clothing, standard flat lay clothing) for training and 1969 pairs for testing, named STGarment. Among the in-the-wild clothing, there are three main displays: clothing worn on a person, clothing laid indoors, and clothing hung on hangers. We use Qwen2-VL-7B [5] to provide prompt annotations for each pair along with 10 discrete attributes (3.5 seconds per image). Additionally, we matched each pair of images with the most structurally similar flat-lay clothing for StructureNet training based on canny image similarity. We employed a clustering approach to filter and construct a memory database for retrieval, which contains 4,000 embedding-landmark pairs from the training set. Please refer to the Appendix for more details.
Method SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ KID $\downarrow$
PBE 0.5278 0.4821 26.55 10.08
IP-Adapter 0.6067 0.4728 14.50 3.747
ReferenceNet 0.6546 0.3979 10.70 1.399
RAGDiffusion 0.6963 0.3684 9.990 1.092
@@
Method SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ KID $\downarrow$
- w/o Retrieval 0.6614 0.3917 10.55 1.354
- w/o StructureNet 0.6576 0.3961 10.65 1.405
- w/o PGEA 0.6926 0.3690 10.22 0.940
Full RAGDiffusion 0.6963 0.3684 9.990 1.092
@@
Implementation details. We initialize the RAGDiffusion with a pre-trained SDXL model and train it on STGarment using an AdamW optimizer with the learning rate of $55e-$ 5 at a resolution of $768\times 768$. The models are trained for $5$ days on $8$ 80GB-A100 GPUs with DeepSpeed [2] ZeRO-2 to reduce memory usage, at a batch size of $15$. $K=4$ in SLLE in Eq. 2. The StructureNet (i.e. embedding encoder) utilizes CLIP-ViT-L/14 image encoder as the backbone and is fine-tuned on nearly 200 million various garment images following DinoV2 [45]. StructureNet is further trained on STGarment with contrastive learning for $4$ days on $4$ 80GB-A100 GPUs at a batch size of $128$. At inference time, we run RAGDiffusion on a single NVIDIA RTX 3090 GPU for $30$ sampling steps with the DDIM sampler [57]. Please refer to the Appendix for more details.
Evaluation protocols. About generation quality, We utilize LPIPS [77] and SSIM [63] to assess the reconstruction accuracy of the generated garments by comparing them with the provided ground truth (GT) images. Additionally, we employ FID [46] and KID [58] metrics to evaluate the realism and authenticity of the generated distributions. In terms of retrieval ability, we evaluate the top-1 accuracy and top-5 accuracy of retrieval results from the memory database $\mathcal{D}$ across 1969 test samples, alongside the average Intersection over Union (IoU) of the corresponding silhouette masks with GT ones. Given the absence of GT masks of test samples in the memory database $\mathcal{D}$, we define the retrieved masks to be accurate if their IoU exceeds $0.85$.
Baselines. Standard garment generation from images is a new task. In general, conditional generation using SD involves three classic methods: IP-Adapter [69], ReferenceNet [28], and ControlNet [75]. ControlNet is primarily designed for spatial-aligned conditional injection, which is not suitable for our task. Therefore, we trained IP-Adapter and ReferenceNet as powerful and intuitive baselines based on SDXL. Additionally, we also trained the classic image-to-image model Paint-by-Example [66] (PBE) as a baseline.
@@ 4.2 Generation comparisons with baselines
Qualitative results. Fig. 3 presents a qualitative comparison between RAGDiffusion and baselines on the STGarment dataset. Paint-by-example has yielded entirely unsuccessful outputs. Meanwhile, IP-Adapter exhibits a high-level perception of garment semantics, generating roughly accurate structure even in layered and side-view situations, although it suffers from severe texture distortion. In contrast, ReferenceNet manages to maintain correct textures in challenging scenarios but encounters structural confusion and distortion. This phenomenon may stem from ReferenceNet’s over-reliance on local texture features, lacking the benefits of a global perspective in challenging cases. RAGDiffusion employs a retrieval-aggregate approach to capture structural information and integrates conditional controls from EP-Adapter, Landmark Guider, and ReferenceNet, yielding promising performance in both structurally faithful and detail-oriented garment conversion. Notably, while all models are trained on the same dataset, RAGDiffusion assimilates high-quality contour landmarks and structure embeddings as external prior to producing visually compelling results that enhance depth and realism.
Figure 5: Ablation study. Without retrieval, RAGDiffusion suffers inaccurate external silhouettes.
Embeddings from StructureNet improve inner structure, while PGEA enhances detail preservation.
Figure 6: Embedding remapping sensitivity to $K$ in Eq. 3.
@@
User study. When metrics like FID and SSIM assess the realism and authenticity of standard garment generation, they tend to be less sensitive to high-frequency information such as contours, collar variations, buttons, and color shifts [16]. We conducted user studies involving 25 participants to evaluate different methods based on user preferences across 200 randomly selected image pairs from the test set. Participants were asked to assign a preference score (ranging from $1\sim 10$) in terms of structural fidelity and detail fidelity for each sample generated by anonymized methods. As illustrated in Fig. 4, we report the distribution of the scores, encompassing medians, means, quartiles, and outliers. Our findings reveal that RAGDiffusion is significantly preferred over all baseline methods in terms of structural fidelity ($mean=8.34$) and detail fidelity ($mean=8.52$). Additionally, the narrow range of outliers in our method indicates a more stable generation across various scenes.
@@ 4.3 Ablation study
Retrieval. One of the key processes of SLLE is to retrieve silhouette landmarks that enhance explicit spatial constraints. Consequently, we removed landmarks as the baseline named ”w/o Retrieval” in Tab. 2 and Fig. 6. The model without retrieval exhibits significant contour errors, particularly around the back collar, sleeve length, and garment length. This is attributed to the fact that retrieval-augmented SLLE incorporates priors from LLM, which transcend visual limitations and aid the generative model in accurate spatial structures of clothing through the use of landmarks. Furthermore, we investigate the role of SLLE in embedding remapping in Sec. 4.4.
@@
StructureNet. The StructureNet, based on contrastive learning, provides latent structure embeddings that are utilized for similarity-based retrieval in SLLE and feature injection in the EP-Adapter. To validate the necessity of our proposed StructureNet, we removed the structure embedding and the subsequent retrieval process as the baseline labeled ”w/o StructureNet”. As illustrated in Tab. 2 and Fig. 6, the model lacking latent structure embedding frequently displays artifacts in the internal structures of the clothing, such as inaccurate collars, missing pockets, and confusion due to occlusions. This demonstrates that structure embedding serves as global guidance and plays a significant role in optimizing the internal structures of apparel.
@@
PGEA. PGEA is employed to mitigate information loss during the VAE encoding and decoding process. As illustrated in Tab. 2 and Fig. 6, the model utilizing PGEA achieves clearer and more accurate edges in high-frequency texture simulation; additionally, it shows significant improvements in the recovery of logos and text. PGEA effectively alleviates the long-standing issue of distortion in fine patterns that has hindered the widespread application of AIGC in commercial scenarios.
@@ 4.4 Component analysis
Latent structure embedding distributions. We further visualize the learned latent embedding distribution in Fig. 7 to gain an in-depth comprehension of how the embeddings work in retrieval and the EP-Adapter. For this purpose, we sample $N=200$ pairs of (in-the-wild cloth embedding $e_{itw}$, standard cloth embedding $e_{std}$) pairs from our dual-tower StructureNet. We then employ t-SNE [61] to project each feature representation into a point within Fig. 7. Notably, 1) the dual-branch embeddings ($e_{itw}$, $e_{std}$) corresponding to the same category exhibit clustering behavior, indicating that the learned priors in StructureNet effectively perform structural alignment cross domains, thus facilitating retrieval based on structural similarity; and 2) representations from different categories are clearly dispersed, demonstrating that the embeddings encompass discriminative structural information, which aids in generation through the EP-Adapter. Lastly, we also visualize the distribution of cosine similarity between a given sample and images from the external memory database $\mathcal{D}$ in Fig. 8. The cosine similarities present a normal distribution, affirming that the constructed embedding library is representative and comprehensive. Furthermore, the examples visualized for different similarity retrieval results effectively illustrate the efficacy of the landmark retrieval mechanism.
Number of retrieved embeddings. As the number $K$ of retrieved nearest neighbors in Eq. 2 during SLLE plays a fundamental role in the properties of the final structure embeddings and landmarks, we demonstrate the fusion sensitivity to $K$ in LABEL:{fig:_knn_curve}. A smaller $K$ value facilitates the sharp edge of landmarks, whereas a larger $K$ value enhances the accurate representation of the reconstructed embedding $e_{itw}^{\prime}$ in Eq. 3. It can be observed that the optimal performance occurs at $K=4$, which also serves as the default parameter setting for other experimental parts.
4.5 Extended applications of RAG
The core concept of RAGDiffusion lies in leveraging retrieval to provide structural determinism, thereby enhancing the quality and fidelity of generated content. This approach has broad applicability beyond garment generation. When equipped with a comprehensive retrieval database and an accurate retrieval mechanism, it can effectively address issues such as incorrect hand structure and prompt control failures, minimizing structural hallucinations.
In extended studies, we employ a training-free RAG system in three tasks in Fig. 9: Pose Control, Finger Refining, and Background Replacing, with retrieval banks consisting of 30 landmarks for each task. As an ablation, we employ SD 1.5 to directly generate images from prompts, which often results in (a) poses that misalign with the descriptions, (b) structural inaccuracies, and (c) outputs of low quality. Subsequently, we utilize GPT-4o to retrieve corresponding landmarks from the external demo banks, employing pre-trained ControlNet [75] or IP-Adapter [69] for conditioning. This intervention significantly improves the quality of the generated outputs and eliminates structural hallucinations. While this represents a training-free toy demonstration, it showcases the extensive potential of RAG-assisted generation, particularly in professional commercial contexts that demand high standards and rigor.
Figure 7: Visualization of embedding distribution with t-SNE. Observe intra-class dual-tower coherence and inter-class divergence.
@@ 4.5 Extended applications of RAG
s Refer to caption
Figure 9: Extended cases of RAG to provide structural determinism, enhancing the quality and fidelity towards human intentions.
@@ 5 Conclusion
In conclusion, RAGDiffusion presents a significant advancement in the generation of standardized clothing assets by effectively addressing the prevalent challenges of structural hallucinations and fidelity in generated images. By integrating Retrieval-Augmented Generation with a dual-process framework—focusing on retrieval-based structure aggregation and omni-level faithful generation—we demonstrate a robust approach to creating high-quality garment images that maintain structural integrity and detail accuracy. Comprehensive experiments and in-depth analysis demonstrate notable performance improvements over existing generative models. We hope that RAGDiffusion will help open avenues for RAG in diverse high-specification, faithful generation tasks, bringing us closer to the goal of professional content creation for all.
@@
In Sec. 6, we demonstrate that RAGDiffusion significantly enhances generalization capabilities by updating the retrieval database without the need for retraining. In Sec. 7, we discuss the limitations of our work and outline directions for future research. Given that standard garment generation is a novel task, we elaborate on its practical significance in Sec. 8. To enhance reproducibility, we provide a detailed description of the network architecture, hyperparameters, and procedural steps in Sec. 9, thoroughly covering the process of RAG system construction. In Sec. 10, we discuss preliminary knowledge relevant to this work. Finally, in Sec. 11, we detail the experimental evaluation process and provide additional results from the ablation study, along with extensive case studies of RAGDiffusion within STGarment.
6 Additional analysis: enhancing generalization through external retrieval databases

### v2

@@ 4.1 Experimental setup
Datasets. We collected 65,131 pairs of (in-the-wild upper clothing, standard flat lay upper clothing) for training and 1969 pairs for testing, named STGarment. Among the in-the-wild clothing, there are three main displays: clothing worn on a person, clothing laid indoors, and clothing hung on hangers. We use Qwen2-VL-7B [2] to provide prompt annotations for each pair along with 10 discrete attributes (3.5 seconds per image). Additionally, we matched each pair of images with the most structurally similar flat-lay clothing for StructureNet training based on canny image similarity. We employed a clustering approach to filter and construct a memory database for retrieval, which contains 4,000 embedding-landmark pairs from the training set. Please refer to the Appendix for more details.
Method SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ KID $\downarrow$ DISTS $\downarrow$
PBE 0.5278 0.4821 26.55 10.08 0.352
ControlNet 0.5225 0.4792 23.02 9.231 0.345
IP-Adapter 0.6067 0.4728 14.50 3.747 0.264
TryOffDiff 0.6124 0.5007 34.34 14.19 0.335
TryOffAnyone 0.6575 0.4335 18.86 7.919 0.245
ReferenceNet 0.6546 0.3979 10.70 1.399 0.224
RAGDiffusion 0.6963 0.3684 9.990 1.092 0.192
@@
Method SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ KID $\downarrow$ DISTS $\downarrow$
Vanilla Model 0.6576 0.3961 10.65 1.405 0.222
+ emb 0.6614 0.3917 10.55 1.354 0.217
+ emb + sil 0.6926 0.3690 10.22 0.940 0.210
Generating mask 0.6426 0.4186 14.55 3.570 0.239
+ emb + sil + PGEA (Full) 0.6963 0.3684 9.990 1.092 0.192
@@
Dataset Viton-HD DC-upper DC-lower DC-dress
Method FID KID FID KID FID KID FID KID
PBE 80.8 52.9 45.0 21.1 158.8 128.3 105.1 74.6
ControlNet 69.9 43.2 48.9 23.6 174.7 161.7 118.3 95.4
IP-Adapter 46.8 17.8 28.8 11.1 111.3 62.7 47.8 19.0
TryOffDiff* 14.0 3.31 38.9 16.7 158.7 130.0 74.3 48.6
TryOffAnyone* 11.4 1.97 24.4 9.30 114.2 79.6 38.4 20.4
ReferenceNet 15.2 3.23 17.8 6.24 124.5 90.2 51.2 22.4
Ours 12.3 1.45 15.9 4.51 40.5 19.4 23.1 6.22
Table 3: Cross dataset evaluation on Viton-HD, DressCode. Model with * means inner-dataset test which is trained on Viton-HD.
Implementation details. We initialize the RAGDiffusion with a pre-trained SDXL model and train it on STGarment using an AdamW optimizer with the learning rate of 5e55e-5 at a resolution of $768\times 768$. The models are trained for 55 days on 88 H20 GPUs with DeepSpeed [31] ZeRO-2 to reduce memory usage, at a batch size of $15$. $K=4$ in SLLE in Eq. 2. The StructureNet (i.e. embedding encoder) utilizes CLIP-ViT-L/14 image encoder as the backbone and is fine-tuned on nearly 2 million various garment images following DinoV2 [34]. StructureNet is further trained on STGarment with contrastive learning for 44 days on 44 H20 GPUs at a batch size of $128$. At inference time, we run RAGDiffusion on a single NVIDIA RTX 3090 GPU for $30$ sampling steps with the DDIM sampler [47]. Please refer to the Appendix for more details.
Evaluation protocols. About generation quality, we utilize LPIPS [63], SSIM [52] to assess the reconstruction accuracy, DISTS [13] to evaluate image similarity on both perceptual and structural level. Additionally, we employ FID [35] and KID [48] metrics to evaluate the realism and authenticity of the generated distributions. In terms of retrieval ability, we evaluate the top-1 accuracy and top-5 accuracy of retrieval results from the memory database $\mathcal{D}$ across 1969 test samples, alongside the average Intersection over Union (IoU) of the corresponding silhouette masks with GT ones. Given the absence of GT masks of test samples in the memory database $\mathcal{D}$, we define the retrieved masks to be accurate if their IoU exceeds $0.85$.
Baselines. We compare RAGDiffusion with recent works TryOffDiff [51] and TryOffAnyone [53] with official checkpoints. We also train 4 classic conditional generation methods on STGarment with SDXL backbone for fair comparison: IP-Adapter [57], ReferenceNet [23], ControlNet [61] and Paint-by-Example [55] (PBE) as baselines.
@@ 4.2 Generation comparisons with baselines
Qualitative results. Fig. 3 presents a qualitative comparison between RAGDiffusion and baselines on the STGarment dataset. TryOffDiff has yielded entirely unsuccessful outputs. Meanwhile, naive ReferenceNet and TryOffAnyone manage to maintain correct textures in challenging scenarios, but encounter structural confusion and distortion. This phenomenon may stem from TryOffDiff and ReferenceNet’s over-reliance on local visual texture features, lacking the benefits of a global perspective in challenging cases. RAGDiffusion employs a retrieval-aggregate approach to capture structural information and integrates conditional controls from EP-Adapter, Landmark Guider and PGEA, yielding promising performance in both structurally faithful and detail-oriented garment conversion. Notably, while trained on the same dataset with naive ReferenceNet, RAGDiffusion assimilates high-quality contour landmarks and structure embeddings as external prior to producing visually compelling results that enhance depth and realism.
Figure 5: Ablation study. Without landmarks, RAGDiffusion suffers inaccurate shape. Embeddings from StructureNet improve inner structure, while PGEA enhances detail preservation.
Figure 6: RAG significantly improves generalization on untrained categories.
@@
User study. When metrics like FID assess the realism and authenticity of standard garment generation, they tend to be less sensitive to high-frequency information such as contours, collar variations, buttons, and color shifts [11]. We conducted user studies involving 100 participants to evaluate different methods based on user preferences across 200 randomly selected image pairs from the test set. Participants were asked to assign a preference score (ranging from $1\sim 10$) in terms of structural fidelity and detail fidelity for each sample generated by anonymized methods. As illustrated in Fig. 4, we report the distribution of the scores, encompassing medians, means, quartiles, and outliers. Our findings reveal that RAGDiffusion is significantly preferred over all baseline methods in terms of structural fidelity ($mean=8.34$) and detail fidelity ($mean=8.52$). Additionally, the narrow range of outliers in our method indicates a more stable generation across various scenes.
@@ 4.3 Ablation study
To validate our core contributions, we define a Vanilla Model that removes structure embedding $\hat{e}_{itw}$, silhouette landmarks $\hat{L}_{sil}$ and PGEA. Specifically, its $\hat{e}_{itw}$ and $\hat{L}_{sil}$ are set to zero vectors while using original SDXL VAE.
@@
Soft guidance of structure embeddings. Since structure embeddings serve as prerequisites for retrieval in SLLE and inputs to EP-Adapter, we first add structure embeddings $\hat{e}_{itw}$ into EP-Adapter of vanilla model (denoted as “+ emb”). As in Tab. 2 and Fig. 6, vanilla model lacking structure embedding frequently exhibits artifacts in clothing internal structures, such as inaccurate collars, missing pockets, and occlusion confusion. This demonstrates that structure embedding provides global guidance and plays a fundamental role in optimizing apparel internal structures.
Hard guidance of silhouette landmarks. The core of RAG is to retrieve silhouette landmarks that enhance explicit spatial constraints. Thus we add landmarks on “+ emb” version as the ablation named “+ emb + sil” in Tab. 2. The model without retrieval exhibits significant contour errors, particularly around the back collar, sleeve length, and garment length. This is attributed to the fact that retrieval-augmented SLLE incorporates priors from LLM, which transcend visual limitations and aid in accurate spatial structures through the use of landmarks. Furthermore, we investigate the role of SLLE in embedding remapping in Sec. 4.5.
Alternative to RAG: generating masks via UNet. As RAG requires a memory database and contrastive learning for embeddings, we try a simpler alternative by employing an SDXL UNet to directly predict landmarks. Specifically, we modify the input of the alternative UNet to a three-channel in-the-wild image $x_{itw}$, with outputs being one-channel predicted landmarks $L_{sil}$ that are directly fed into RAGDiffusion’s denoising UNet. However, this alternative produces landmarks with noticeable errors and degraded generation quality, due to limitations in visual understanding and the absence of proper pre-trained weights.
@@
PGEA. PGEA is employed to mitigate information loss during the VAE encoding and decoding process. We add PGEA on “+ emb + sil” version as Full Model. The full model utilizing PGEA achieves clearer and more accurate edges in high-frequency patterns simulation; additionally, it shows significant improvements in the recovery of logos and text. PGEA greatly alleviates the long-standing issue of distortion in detail and makes RAGDiffusion commercially viable, the first in industrial garment restoration.
4.4 Additional rationale for introducing RAG
In addition to structural faithfulness, RAGDiffusion demonstrates excellent generalizability and robustness across untrained categories and scenes simply by updating retrieval database, without the need for collecting expensive paired data and retraining. Besides, please refer to Appendix for another rationale of human-interpretable control.
Cross category evaluation. RAGDiffusion is trained on upper-body clothing data and has not encountered lower-body garments during training. To validate the generalizability, we collect $856$ flat-lay lower-body/dress images (no in-the-wild image is needed) and incorporate them with silhouette masks into the external memory database through StructureNet. Subsequently, we test it on $50$ lower-body in-the-wild clothing samples, using a ReferenceNet as baseline. The results in Fig. 6 demonstrate that retrieval significantly improves generalization capabilities, serving as a cost-effective zero-shot generalizing solution. The quantitative results on lower-body/dress categories in Tab. 3 also validate the generalizability boost.
Cross dataset evaluation. Despite zero exposure to the Viton-HD [12] and DressCode [21] datasets during training, RAGDiffusion shows strong out-of-distribution (OOD) compatibility over baselines without tuning illustrated in  Tab. 3. This boost in generalization increases the operational maturity of RAGDiffusion, enabling it to effectively handle various OOD images submitted by users.
@@ 4.5 Component analysis
Latent structure embedding distributions. We further visualize the learned latent embedding distribution in Fig. 7(b) to gain an in-depth comprehension of how the embeddings work in retrieval and the EP-Adapter. For this purpose, we sample $N=200$ pairs of (in-the-wild cloth embedding $e_{itw}$, standard cloth embedding $e_{std}$) pairs from our dual-tower StructureNet. We then employ t-SNE [50] to project each feature representation into a point within Fig. 7. Notably, 1) the dual-branch embeddings ($e_{itw}$, $e_{std}$) corresponding to the same category exhibit clustering behavior, indicating that the learned priors in StructureNet effectively perform structural alignment cross domains, thus facilitating retrieval based on structural similarity; and 2) representations from different categories are clearly dispersed, demonstrating that the embeddings encompass discriminative structural information, which aids in generation through the EP-Adapter. Lastly, we also visualize the distribution of cosine similarity between a given sample and images from the external memory database $\mathcal{D}$ in Fig. 8. The cosine similarities present a normal distribution, affirming that the constructed embedding library is representative and comprehensive. Furthermore, the examples visualized for different similarity retrieval results effectively illustrate the efficacy of the landmark retrieval mechanism.
Number of retrieved embeddings. As the number $K$ of retrieved nearest neighbors in Eq. 2 during SLLE plays a fundamental role in the properties of the final structure embeddings and landmarks, we demonstrate the fusion sensitivity to $K$ in LABEL:{fig:_tsne}(a). A smaller $K$ value facilitates the sharp edge of landmarks, whereas a larger $K$ value enhances the accurate representation of the reconstructed embedding $e_{itw}^{\prime}$ in Eq. 3. It can be observed that the optimal performance occurs at $K=4$, which also serves as the default parameter setting for other experimental parts.
Figure 7: Illustrations of component analysis.
@@ 5 Conclusion
In conclusion, RAGDiffusion presents a significant advancement in the generation of standardized clothing assets by effectively addressing the prevalent challenges of structural hallucinations and texture fidelity. By integrating retrieval-based structure aggregation to get soft and hard guidance, as well as omni-level faithful generation pipeline, RAGDiffusion is capable of synthesizing both structure and texture faithful clothing assets. Comprehensive experiments and in-depth analysis demonstrate notable performance improvements over existing models. We hope that RAGDiffusion helps open avenues for RAG in diverse high-specification faithful generation tasks, bringing us closer to the goal of professional content creation for all.
Acknowledgements
This work is supported by the Science and Technology Commission of Shanghai Municipality under research grant No. 25ZR1401187.
[…]
@@ Acknowledgements
In Sec. 6, we demonstrate two additional benefits brought by the RAG system: enhanced generalizability, and human-interpretable control through landmark manipulation. In Sec. 7, we discuss the limitations of our work and outline directions for future research. Given that standard garment generation is a novel task, we elaborate on its practical significance in Sec. 8. To enhance reproducibility, we provide a detailed description of the network architecture, hyperparameters, and procedural steps in Sec. 9, thoroughly covering the process of RAG system construction, including dataset. In Sec. 10, we discuss preliminary knowledge relevant to this work. Finally, in Sec. 11, we detail the experimental evaluation process and provide additional results from the ablation study, along with extensive cases of RAGDiffusion within STGarment, Viton-HD and DressCode.
Figure 9: Generalization comparison on crazy difficult garments.
Figure 11: Users can modify the final visual presentation by replacing recommended silhouette masks/landmarks with one of $K$ candidates before UNet denoising, highlighting RAGDiffusion’s advantages in human-interpretable control and retrieval-based recommendation manipulation.
6 Additional rationale for introducing RAG
In addition to enhancing structural determinacy and eliminating structural distortion through the assimilation of structural landmarks and external knowledge, we have also discovered two additional benefits brought by the RAG system: enhanced generalizability, and human-interpretable control through landmark manipulation.
6.1 Enhanced generalizability due to RAG